OpenStudy (anonymous):

You are given a vector in the xy plane that has a magnitude of 90.0 units and, a y component of -41.0 units. Assuming the x component is known to be positive, specify the vector, V, if you add it to the original one, would give a resultant vector that is 89.0 units long and points entirely in the -x direction.

OpenStudy (anonymous):

Okay so the previous question to the one I'm confused on asks me to find two possible answers for this vector's x component which I got -80.1,80.1 using the pytheorem's therom. For this particular question, if I'm being asked to get a direction that points in the -x direction how is that going to be possible if the x component is assumed positive?

OpenStudy (amistre64):

You are given a vector in the xy plane that has a magnitude of 90.0 units and, a y component of -41.0 units. Assuming the x component is known to be positive, specify the vector, V, if you add it to the original one, would give a resultant vector that is 89.0 units long and points entirely in the -x direction. <-89,0>

OpenStudy (amistre64):



OpenStudy (anonymous):

So that vector V would be -89?

OpenStudy (amistre64):

V would be what we add to the first part to get to <-89,0>

OpenStudy (amistre64):

<a,-41> +<x , y > ----------- <-89, 0> y = 41 in this case; we just need to determine a and x we can do a from the pythag thrm; we know: 90^2 = 41^2 + a^2

OpenStudy (amistre64):

a^2 = 90^2 - 41^2 a = sqrt(90^2-41^2) x+a = -89, so x = -89-a x = -89 - sqrt(90^2-41^2)

OpenStudy (amistre64):

gotta make sure i did that right :)

OpenStudy (amistre64):

<-169.12 , 41> might be our V vector

OpenStudy (amistre64):

doesnt make sense to me for some reason

OpenStudy (amistre64):

unless this formula is reading a over in quadrant 4 instead of quadrant 3

OpenStudy (amistre64):

<-8.88, 41> looks better from my drawing

OpenStudy (amistre64):